How many liters of water should be evaporated from 60 liters of a 28% salt solution so that the remaining solution contains 32% salt?

Answer by stanbon(75887)

How many liters of water should be evaporated from 60 liters of a 28% salt solution so that the remaining solution contains 32% salt?
---------
Equation:
water - water = water
0.72*60 - x = 0.68(60-x)
-------------------------------
72*60 - 100x = 68*60 - 68x
------
-32x = -4*60
-----
x = (1/8)60
x = 7.5 liters (amt of water to evaporate)
-----------
Cheers,
Stan H.